ABC Company, a manufacturer of automobile tires claims that the average life of the product is 45,600 miles. A random sample of 15 tires was chosen and resulted in a mean life of 43,500 miles with a standard deviation of 3,000 miles. Based on the above results, can you accept the manufacturer’s claim? Use α = 0.05



1
Expert's answer
2021-12-13T15:45:50-0500

"n=15 \\

\bar{x} = 43500 \\

s = 3000 \\

H_0: \mu = 45600 \\

H_1: \mu < 45600 \\

α = 0.05"

Test-statistic

"t = \frac{\bar{x} - \mu}{s / \sqrt{n}} \\

t = \frac{43500 -45600}{3000 / \sqrt{15}} \\

t = -2.711 \\

d.f. = n-1 = 15-1 = 14 \\

t_{14, 0.05} = -1.761"

Since "t < t_{14, 0.05}" we reject H0 at 5 % level of significance and conclude that the company’s average product’s life is less than 45600 miles.